OpenStudy (anonymous):

[9.05] Solve 3x2 + 4x + 2 = 0. Round solutions to the nearest hundredth.

OpenStudy (lgbasallote):

use quadratic formula

OpenStudy (lgbasallote):

need more help?

OpenStudy (anonymous):

Yeah

OpenStudy (lgbasallote):

quadratic formula: \[\LARGE x = \frac{-b \pm \sqrt{b^2 = 4ac}}{2a}\]

OpenStudy (lgbasallote):

that = is a -

OpenStudy (anonymous):

\[\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array} \] you are given\[ 3x^2 + 4x + 2 = 0\] can you tell me what a b and c are?

OpenStudy (lgbasallote):

so substitute a = 3 b = 4 c = 2

OpenStudy (fellowroot):